OpenStudy (anonymous):

On the critical portion of the SAT, the raw score is calculated as follows: 1 point is awarded for each correct answer, and 1/4 point is deducted for each wrong answer. If Ellen answer all q questions on the test and earned a raw score of 10, how many questions did she answer correctly?<-- LOL @ the question Answer choices: (A) q-10 (B) q/5 (C) q/5-10 (D) q-10/g (E) 8+q/5

OpenStudy (ganpat):

q = total attempt x = right attempt so, for right answer 1 mark and wrong answer, -1/4 marks.. 1(x) - 1/4 (q-x) = 10.. can u solve for x ??

OpenStudy (anonymous):

i think so. hold on a second

OpenStudy (anonymous):

x=q/5+8

OpenStudy (anonymous):

right?

OpenStudy (ganpat):

1(x) - 1/4 (q-x) = 10 4x -q + x = 40 5x = 40 + q x = 8 + q/5... perfect ... :) So option 5...

OpenStudy (anonymous):

yay! E.. Thank you so much!

OpenStudy (ganpat):

anytime... :)